OpenStudy (lxelle):

Solve -3e^(-3x) tanx + e^(-3x) sec^2 (x) = 0

OpenStudy (lxelle):

So, e^(-3x) (sec^2 (x) -3 sin(x)) = 0 (1+ tan^2 (x) - 3sin (x)) = 0

OpenStudy (lxelle):

@phi

OpenStudy (anonymous):

Why did \(\tan x\) suddenly become \(\sin x\)?

OpenStudy (lxelle):

mistake. thanks for pointing out